1.76 Create a data set. Create a data set with seven observations for which the median would change by a large amount if the smallest observation were deleted.
Answer #1

Dataset; {0, 1, 2, 3, 32, 33, 34}

In this dataset, median = (n+1)/2th term

= (7 + 1)/2 th term

= 4th term

= 3

When the smallest observation is deleted, dataset is {1, 2, 3, 32, 33, 34}

Median = (6+1)/2th term

= 3.5th term

= (3rd term + 4th term)/2

= (3 + 32)/2

= 17.5

The median have increase multiple folds when the smallest observation was deleted.

The